Calculate the degree of hydrolysis of a decinormal KCN solution at 25 degree celcius. The dissociation constant of HCN is 7.2 x 10 - 10 and ionic product of water is 10 – 14 . Find the pH of the solution.

  1. Calculate the degree of hydrolysis of a decinormal KCN solution at 25 degree celcius. The dissociation constant of HCN is 7.2 x 10 - 10  and ionic product of water is 10 – 14 .
  2. Find the pH of the solution.

Grade:11

1 Answers

Vikas TU
7 years ago
The reaction for the dissociation of KCN into the solution is:
KCN  + H2O ====>         H+         +           CN-        +         OH-
[H+] = [OH-]   +   [CN-]
[H+]^2 = Ka[HCN] + 10^-14
          = 7.2*10^-10*10^-1  + 10^-14
          = 72*10^-12 + 0.01*10^-12
[H+] = 72*10^-6 M
pH = -log(72*10^-6) = 4.14